OpenStudy (anonymous):

You are arguing over a cell phone while trailing an unmarked police car by 25 m; both your car and the police car are traveling at 110 km/h. Your argument diverts your attention from the police car for 2.0 s (long enough for you to look at the phone and yell, “I won't do that!”). At the beginning of that 2.0 s, the police officer begins braking suddenly at 5.0 m/s^2. If you too brake at 5.0 m/s^2, what is your speed when you hit the police car?

OpenStudy (jamesj):

Nice. For both cars, you want to write down an equation for the position of each car at time t. At time t = 0, let us say that position of police car, write s_p, s_p(t=0) = 25 At time t = 0, that means the position of your car, write s, s(t=0) = 0 i.e., s_p(0) = 25 s(0) = 0 Now both cars are traveling at an initial speed of 110 km/hr. That is u = 30.6 m/s If neither cars decelerated, their equations of motion would be s_p(t) = 25 + 30.6t s(t) = 30.6t Now you figure out how to incorporate the deceleration. The two cars collide when s_p(t) = s(t)

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

wow why so many ppl looking at my question lol

OpenStudy (anonymous):

this problem requires only the concept of uniform acceleration u have to understand the question well thats because when more than 2 look at a question,attention is automatically grabbed!

OpenStudy (anonymous):

haha ah k